Identify five differences between financial and managerial accounting.

If you investing in a business, which of the three types of financial statements you would want to review. Why?

A company shows the following selected financial information from activities for the current year.

Gross sales $225,000

Current assets $40,000

Long-term assets $100,000

Accounts Payable $16,000

5 Year Note Payable $44,000

Net Income $7,200

Outstanding shares 5,000

Par value of shares $9 per share

Retained Earnings $35,000 (includes current net income)

Calculate the following from the above information provided:

Current ratio

Debt-to-equity ratio

Return on Owner’s equity

Earnings per share

Explain how financial ratios help managers monitor their own efficiency and effectiveness.

Support your position/opinions with reference to your readings. Your posting should be about 300 words and include a web link to your online article, video, blog, etc.
Answer #1
Point of difference Financial accounting Managerial accounting
Aggregation Financial accounting reports on the results of an entire business. Managerial accounting almost always reports at a more detailed level, such as profits by product etc.
Efficiency Financial accounting reports on the profitability of a business. Managerial accounting reports on specifically what is causing problems and how to fix them.
Reporting focus Financial accounting is oriented toward the creation of financial statement. Managerial accounting is more concerned with operational reports.
Time period It is concerned with the financial results that a business has already achieved, so it has a historical orientation. It may address budgets and forecasts, and so can have a future orientation.
Valuation It is addresses the proper valuation of assets and liabilities. whereas, it is concerned with only productivity.
The three major financial statements reviewed at the time investment in business are :-
1) Balance sheet (2) Income statement (3) Cash flow statement
All three accounting statements are important for understanding and analyzing a company's performance from multiple angles. The income statement provides deep insight into the core operating activities that generate earnings for the firm. The balance sheet and cash flow statement, however, focus more on the capital management of the firm in terms of both assets and structure.
Financial ratios
Current ratio = Current asset/ current liability
40000/16000 =2.5
Debt to equity ratio = Total Liabilities/Total shareholder's equity
160000/80000 = 2
Return on Onwer's equity = Net income/ Shareholder's equity
7200/35000 = 20.57%
Earnings per share = Net income/ shares outstanding
7200/5000 = 1.44 per share
The financial ratios for efficiency and effectiveness assess a company's operations and profitability. Financial ratios, including efficiency and effectiveness ratios, are based on income statement and balance sheet items.
The three main efficiency ratios are days sales outstanding, inventory turnover ratio and accounts payable-to-sales ratio.
Effectiveness ratios include return on sales, return on assets and return on equity. They indicate how effective management has been in using shareholders' equity and company assets to generate an acceptable rate of return.